# DeepExtractor: Time-domain reconstruction of signals and glitches in gravitational wave data with deep learning

### v1

@@ Abstract
Gravitational wave (GW) interferometers, such as LIGO, Virgo, and KAGRA, detect faint signals from distant astrophysical events, such as binary black hole mergers. However, their high sensitivity also makes them susceptible to background noise, which can obscure these signals. This noise often includes transient artifacts called “glitches”, that can mimic genuine astrophysical signals or mask their true characteristics. Fast and accurate reconstruction of both signals and glitches is crucial for reliable scientific inference. Accurate GW signal reconstruction facilitates parameter estimation of the GW source, while glitch reconstruction—challenging due to the unmodeled nature of glitches—supports glitch mitigation and enables mock data challenges involving realistic noise artifacts. In this study, we present DeepExtractor, a deep learning framework that is designed to reconstruct signals and glitches with power exceeding interferometer noise, regardless of their source. We design DeepExtractor to model the inherent noise distribution of GW interferometers, following conventional assumptions that the noise is Gaussian and stationary over short time scales. It operates by predicting and subsequently subtracting the noise component of the input mixture of noise and signal or glitch, retaining only the clean reconstruction. Our innovative approach achieves superior generalization capabilities for arbitrary signals and glitches compared to methods that directly map mixture inputs to the clean training waveforms, yielding precise reconstruction of time-domain signals and glitches. We focus on applications related to glitches and validate DeepExtractor’s effectiveness through three experiments: (1) reconstructing simulated glitches injected into detector noise, (2) comparing its performance with the state-of-the-art BayesWave algorithm, and (3) analyzing real data from the Gravity Spy dataset to demonstrate effective glitch subtraction from LIGO strain data. Our proposed model achieves a median mismatch of only $0.9\%$ for simulated glitches, outperforming several deep learning baselines. Additionally, DeepExtractor surpasses BayesWave in glitch recovery, offering a dramatic computational speedup by reconstructing one glitch sample in approximately $0.1$ seconds on a CPU, compared to BayesWave’s processing time of approximately one hour per glitch.
@@ I Introduction
The first detection of gravitational waves (GWs) from a binary black hole (BBH) merger in 2015 ushered in a new era of astrophysics [1]. Since then, Advanced LIGO [2] and Advanced Virgo [3] detectors have made confident detections of $\sim 90$ compact binary coalescence (CBC) GW events [4, 5, 6], accumulated over three observing runs (O1, O2, O3). With the fourth observing (O4) run currently underway, hundreds of more detections are expected from the enhanced sensitivity of GW detectors [7] and the introduction of KAGRA [8], Japan’s underground detector.
@@
Model-free algorithms, Omicron [17] and Gravity Spy [18], are utilized for glitch detection and characterization, respectively. These tools visualize and identify glitches by analyzing excess power in Q-scans (spectrograms) [19], which are time-frequency representations. Glitches exhibit a broad range of time-frequency morphologies, which Gravity Spy classifies into over 20 distinct categories. Figure 1 illustrates a representative glitch from the ‘tomte’ class, showing its time-series data alongside the corresponding Q-scan used in Gravity Spy analysis.
Glitches can resemble astrophysical signals, increasing false positives  [20, 21, 22, 23, 24, 25, 26, 27, 28] and hindering subsequent analyses. However, when glitches overlap with a GW signal they can significantly bias parameter estimation [29, 30, 31] and must be carefully modeled and removed before the GW signal can be used to make any measurements. This is a commonplace occurrence, with $\sim 25$ out of the $\sim 90$ confident detections so far requiring some form of glitch mitigation [6, 29, 32, 33, 34, 35].
@@
(a) Time series
(b) Q-scan
Figure 1: An example of a ‘tomte’ glitch in LIGO Livingston at GPS time $1263921313$ with an signal-to-noise ratio (SNR) of $48.6$.
@@
DeepExtractor demonstrates superior accuracy in reconstructing simulated glitches and generalizes well to unseen distributions, outperforming several deep learning models. We also present a comparison of glitch reconstructions by DeepExtractor to those obtained using the state-of-the-art BayesWave [41], an advanced algorithm used in GW data analysis to model and subtract glitches from the strain data. Finally, we showcase DeepExtractor’s effectiveness in reconstructing real Gravity Spy [18] glitch events from LIGO’s third observing run, verified through Q-scans and time series plots of the detector data before and after glitch subtraction. This marks the first comprehensive application of deep learning for such tasks in GW data analysis.
This paper is structured as follows; in section II, we discuss the state-of-the-art in glitch modelling and mitigation in LIGO and Virgo and state-of-the-art machine learning research in time series denoising. In section III, we discuss our method, which includes an overview of our framework, our model architecture, the generation of training data, and the experiments implemented to show our DeepExtractor’s efficacy. In section IV, we present the results of the experiments and section V discusses the conclusions of our research.
@@ II.1 GW glitch mitigation
A variety of methods have been developed to model and subtract glitches from the strain data that leverage machine learning [42] and parametric [43, 44] models. However, up to O3, BayesWave [41, 45, 29, 46] and gwsubtract [47, 25] [48] have been the only algorithms used for modeling and subtracting glitches in LIGO-Virgo-KAGRA data analyses [48]. While gwsubtract relies on strongly correlated auxiliary witness sensors to linearly subtract glitches from the strain data, BayesWave models glitches using the detector strain only.
BayesWave models the detector strain $h(t)$ in each interferometer as a linear combination:
$h(t)=n(t)+s(t)+g(t),$ (1)
where $n(t)$ represents Gaussian noise, $s(t)$ an astrophysical signal, and $g(t)$ a glitch. Non-Gaussian components, including signals and glitches, are represented as sums of sine-Gaussian (Morlet-Gabor) wavelets:
$\Psi(t;\vec{\theta})=Ae^{-(t-t_{0})^{2}/\tau^{2}}\cos\left(2\pi f_{0}(t-t_{0})+\phi_{0}\right),$ (2)
where $\vec{\theta}=\{t_{0},f_{0},Q,A,\phi_{0}\}$ denotes the wavelet parameters: central time ($t_{0}$), central frequency ($f_{0}$), quality factor ($Q$), amplitude ($A$), and phase offset ($\phi_{0}$) with $\tau=Q/2\pi f_{0}$. BayesWave performs Bayesian inference to estimate the wavelet parameters and the total number of wavelets of a glitch using a Reversible-Jump Markov Chain Monte Carlo (RJMCMC) sampler, which allows for the flexibility of models with variable dimensions.
@@
II.2 Deep learning for signal denoising
Reconstructing GWs and glitches parallels time-series denoising tasks in fields like speech enhancement, acoustic processing, and audio source separation, all of which have seen significant advancements. Traditional methods such as dictionary learning, wavelet shrinkage, and singular value decomposition (SVD) have been effective [50, 51, 52]. However, supervised deep learning has revolutionized performance, particularly in data-rich applications like image denoising and speech enhancement [53, 54].
Time vs. time-frequency domains: Networks for denoising can process data in the time or time-frequency domain, each with trade-offs [55]. Time-domain approaches analyze raw waveforms directly, leveraging deep learning’s ability to extract features automatically [56, 57, 58, 59, 60]. Conversely, time-frequency representations like the STFT offer detailed signal characterizations, enhancing learning [61, 62, 54, 63]. However, STFT-based methods require pre- and post-processing steps, increasing computational load [64]. Key STFT parameters, such as frame size and overlap, affect time and frequency resolution but are often fixed rather than optimized jointly with network parameters.
Phase considerations in STFT: Most STFT-based denoising models use only the magnitude spectrogram, assuming the mixture (input) and source (target) phases are equivalent [65]. While effective, this assumption is not always valid, especially in signals with overlapping harmonics and multiple studies have shown the importance of phase for denoising performance [66, 67]. Iterative methods like Griffin-Lim [68] can approximate the source phase but are computationally intensive and may yield suboptimal results [69]. Recent approaches explore complex spectrograms that preserve phase or estimate it directly, showing improved performance [70, 71, 72].
Deep learning architectures for denoising: Popular architectures for denoising include convolutional neural networks (CNNs), convolutional autoencoders, and U-Nets. CNNs have been applied to tasks such as electrocardiogram denoising, speech enhancement, image denoising and source separation in time-frequency domains [73, 74, 53, 75]. Convolutional autoencoders compress signals into a latent representation and reconstruct them, achieving success in areas such as speech enhancement, health management and fault diagnosis [76, 77, 78, 79, 80]. However, deep autoencoders risk information loss in the compressed layer, especially with increasing the number of layers [81].
U-Nets [40] mitigate this issue with skip connections that retain features across layers, achieving superior results in tasks like acoustic source separation and image denoising [82, 58, 83]. Their ability to combine feature preservation with a large receptive field has made them a state-of-the-art choice for signal denoising [84].
Figure 2: An overview of DeepExtractor’s reconstruction approach. Two seconds of detector strain $h(t)$ is processed into magnitude and phase spectrogramns using an STFT. This is fed through our network, which maps the instance to the magnitude and phase spectrograms of the underlying background noise. We then use an inverse STFT to yield to corresponding time series for $n(t)$. An estimation of the underlying signal or glitch can be simply calculated as $h(t)-n(t)$.
II.3 Deep denoisers for GW data analysis
@@
U-Net models have been applied to denoise BBH signals using time-frequency representations [90]. Adapted from seismic signal separation studies [91], this approach leverages the real and imaginary parts of spectrograms to predict separate masks for signal and noise. These masks, applied to input spectrograms, allow reconstruction of the time-domain signal via inverse STFT, yielding strong performance on real astrophysical events.
@@ III.1.1 Overview of framework
DeepExtractor is a deep learning framework designed to reconstruct signals or glitches in GW data with power exceeding the underlying detector noise. An overview of the framework is illustrated in Figure 2. Built using PyTorch’s deep learning library [93], DeepExtractor’s codebase is publicly accessible on GitLab [^1] https://git.ligo.org/tom.dooney/deepextractor.git. DeepExtractor is agnostic to the origin and morphology of excess power in the detector, whether it is a signal or a glitch. Throughout this paper, we use $g(t)$ to denote either a signal or a glitch for convenience, given our primary focus on glitch-related applications. Although we often refer to events as glitches, this is solely for convenience, as the model does not distinguish between the origins of excess power.
The framework operates under the first two models of BayesWave described in Section II.1, where the data contains either an astrophysical signal or a glitch. These models correspond to BayesWave’s signal-only or glitch-only modes. Currently, DeepExtractor does not handle the third model assumption, where a glitch overlaps with an astrophysical signal (BayesWave’s signal+glitch mode). While this capability is critical for a fully-functional glitch mitigation algorithm, this study focuses on the simpler signal-only and glitch-only cases as a foundational step. Potential approaches for addressing the signal+glitch scenario are discussed in Section V.
We follow BayesWave’s assumption in Eq. 1 that a glitch $g(t)$ interacts additively with the strain data in a real detector setting. Our goal is to map the input mixture $h(t)$ to the noise component $n(t)$, where $h(t)$ is represented as:
@@ III.1.2 Training strategy
The DeepExtractor framework utilizes a neural network that directly outputs the background noise component from the data in the spectrogram domain. Our core idea is that the background noise is easier to model by our network because it is simpler and more stationary than the diverse glitches.
@@
DeepExtractor is trained to learn the mapping in the time-frequency domain using invertible STFT transforms. Modeling both magnitude and phase is essential for effective signal and glitch reconstruction in the time domain. To provide the inputs and targets, we first transform the time-domain signals $h(t)$ and $n(t)$ into the time-frequency domain by computing their complex spectrograms, $h(t,f)$ and $n(t,f)$, using the STFT (see Section III.2.1). These spectrograms are then decomposed into magnitude and phase components, which are concatenated channel-wise before being fed into the network. Thus, DeepExtractor takes the magnitude and phase components of $h(t,f)$ as input and outputs the corresponding components of the estimated noise $\hat{n}(t,f)$.
@@ III.1.3 U-Net Architecture
Figure 3: The DeepExtractor architecture implemented on STFT data. The two input and output channels process the magnitude and phase components of the STFT respectively.
Motivated by the discussion in Section II, DeepExtractor adopts a U-Net architecture for its ability to capture both high- and low-level features through skip connections. Skip connections concatenate encoder features with their corresponding decoder features, as illustrated in Figure 3. These connections preserve fine-grained details while enabling a large receptive field, which is particularly advantageous for analyzing spectrogram representations of signals or glitches with broad time-frequency ranges. U-Nets effectively reconstruct inputs without significant information loss.
DeepExtractor is a 2D model designed to process magnitude and phase spectrograms. It accepts two input channels and outputs two spectrograms of the same dimensionality, as illustrated in Figure 3.
@@
Time-Domain Variant: We also experiment using a 1D time-domain counterpart of DeepExtractor’s U-Net architecture (UNET1D). This replaces all 2D operations with 1D equivalents and processes a single input/output channel. Specifically, the 1D model maps the raw, whitened time series $h(t)$ to the noise component $n(t)$, omitting the STFT and iSTFT transformations used in the 2D variant, as shown in Figure 2.
@@ III.2.1 Simulated data
To properly validate our model, we first implement a fully simulated approach for both training and validation, simulating the background noise and glitches. We simulate white noise training targets $n(t)$ with a duration of 2 seconds at a sampling rate of $4,096\,$ Hz, yielding 8,092 data points. The simulated noise has a mean of 0 and variance of $f_{s}/2=2,048\,$Hz, where $f_{s}$ is the sampling rate. We generate 250,000 background samples for training and 25,000 for validation.
To create the corresponding input samples, we inject various analytical waveforms, represented by $g(t)$, into the background noise $n(t)$, as per Eq. 3. These waveforms are drawn from five proxy glitch classes: chirp, sine, sine-Gaussian, Gaussian pulse, and ringdown, as shown in Figure 4. Each waveform class has its own parameters that are randomly sampled during generation, such as frequency range (from 111\,1 Hz to the Nyquist frequency of $2,048\,$Hz), bandwidth, and duration (randomly chosen between 0.1250.1250.125\,0.125s and 222\,2s). For more information on these training classes, please see Appendix VII.1 and our code repository. The SNR is varied between 1 and 250, scaled according to
@@
where $g(f)$ and $S_{n}(f)$ represent the Fourier transform of the injected waveform and the detector noise PSD, respectively [97]. The scaling ensures that each waveform achieves the desired SNR within the specified range.
To improve generalization and avoid overfitting, we inject linear combinations of glitches into each background sample, instead of using a single glitch $g(t)$, where:
@@
and $k\in[1,30]$. Thus, $\mathbf{g}$ will replace $g(t)$ in Eq. 3 when describing the training data moving forward.
This allows for between 1 and 30 glitches per sample, sampled uniformly across the five classes. This follows BayesWave’s method of modeling signals and glitches as sums of sine-Gaussian wavelets (Eq. 6), but increases the space of analytical models with additional classes to improve the network’s generalization ability.
Additionally, glitches can be scaled to low SNRs, so some training samples include injections with very little contribution to the data. This approach helps the model to indirectly learn the background PSD, enhancing its capability to distinguish and reconstruct subtle glitch features while maintaining robustness across diverse detector noise characteristics.
To prepare the data for training DeepExtractor, we scale (standardize) the inputs and outputs using StandardScaler from sklearn [98], which transforms the data so that it has a mean of 0 and a standard deviation of 1. This standardization method assumes a Gaussian distribution of the data and is appropriate given the Gaussian-like characteristics of the background noise. We fit the scaler to the model input only where glitches are present, only transforming the background noise target with the input-fitted scaler. This ensures that the input and output occupy the same standardized space. This preprocessed data is used to train the time-domain benchmarks for DeepExtractor.
Short time Fourier transform (STFT): As mentioned in Section III.1.2, DeepExtractor and the 2D benchmarks are trained on spectrograms calculated from the above preprocessed data using the STFT. Since the time-domain data is already standardized, no additional scaling is performed on the spectrograms. Figure 5 shows examples of the magnitude and phase components for both the input and target spectrograms. We choose STFT parameters to ensure compliance with the Constant Overlap-Add (COLA) condition [99] (please see Appendix VII.2 for more information). A Hann window is employed to minimize spectral leakage, further improving the quality of the spectrogram representation [100]. These chosen settings yield complex spectrograms with dimensions of $257\times 257$, which we separate into magnitude and phase spectrograms. To investigate the effect of varying time and frequency resolutions on signal and glitch reconstruction, we also experiment with alternate parameter configurations that produce spectrograms of dimensions $129\times 129$ and $65\times 129$ while ensuring that all configurations satisfy the COLA condition.

### v3

# Time-domain reconstruction of signals and glitches in gravitational wave data with deep learning

@@ Abstract
Gravitational wave (GW) detectors, such as LIGO, Virgo, and KAGRA, detect faint signals from distant astrophysical events. However, their high sensitivity also makes them susceptible to background noise, which can obscure these signals. This noise often includes transient artifacts called ‘glitches’, that can mimic genuine astrophysical signals or mask their true characteristics. In this study, we present DeepExtractor, a deep learning framework that is designed to reconstruct signals and glitches with power exceeding interferometer noise, regardless of their source. We design DeepExtractor to model the inherent noise distribution of GW detectors, following conventional assumptions that the noise is Gaussian and stationary over short time scales. It operates by predicting and subtracting the noise component of the data, retaining only the clean reconstruction of signal or glitch. We focus on applications related to glitches and validate DeepExtractor’s effectiveness through three experiments: (1) reconstructing simulated glitches injected into simulated detector noise, (2) comparing its performance with the state-of-the-art BayesWave algorithm, and (3) analyzing real data from the Gravity Spy dataset to demonstrate effective glitch subtraction from LIGO strain data. We further demonstrate its potential by reconstructing three real GW events from LIGO’s third observing run, without being trained on GW waveforms. Our proposed model achieves a median mismatch of only $0.9\%$ for simulated glitches, outperforming several deep learning baselines. Additionally, DeepExtractor surpasses BayesWave in glitch recovery, offering a dramatic computational speedup by reconstructing one glitch sample in approximately $0.1$ seconds on a CPU, compared to BayesWave’s processing time of approximately one hour per glitch.
@@ I Introduction
The first detection of gravitational waves (GWs) from a binary black hole (BBH) merger in 2015 ushered in a new era of astrophysics [1]. Since then, Advanced LIGO [2] and Advanced Virgo [3] detectors have made confident detections of $\sim 90$ compact binary coalescence (CBC) GW events [4, 5, 6, 7, 8, 9, 10, 11, 12], accumulated over three observing runs (O1, O2, O3). With the fourth observing (O4) run currently underway, hundreds of more detections are expected from the enhanced sensitivity of GW detectors [13].
@@
Model-free algorithms, Omicron [22] and Gravity Spy [23], are utilized for glitch detection and characterization, respectively. These tools visualize and identify glitches by analyzing excess power in Q-scans (spectrograms) [24], which are time-frequency representations. Glitches exhibit a broad range of time-frequency morphologies, which Gravity Spy classifies into over 20 distinct categories.
Certain glitches can resemble astrophysical signals (eg. blip glitches have similar time-frequency morphology to high-mass BBH signals), increasing false positives  [25, 26, 27, 28, 29, 30, 31, 32, 33] and hindering subsequent analyses. However, when glitches overlap with a GW signal they can significantly bias parameter estimation [34, 35, 36] and must be carefully modeled and removed before the GW signal can be used to make any measurements. This is a commonplace occurrence, with $\sim 25$ out of the $\sim 90$ confident detections so far requiring some form of glitch mitigation [6, 34, 37, 38, 39, 40].
@@
DeepExtractor demonstrates superior accuracy in reconstructing simulated glitches and generalizes well to unseen distributions, outperforming several deep learning models. We also present a comparison of glitch reconstructions by DeepExtractor to those obtained using the state-of-the-art BayesWave [46], an advanced algorithm used in GW data analysis to model and subtract glitches from the strain data. We further demonstrate DeepExtractor’s effectiveness in reconstructing real Gravity Spy [23] glitches from LIGO’s third observing run, verified through Q-scans and time series plots of the detector data before and after glitch subtraction. This marks the first comprehensive application of deep learning for such tasks in GW data analysis. Finally, we use DeepExtractor to reconstruct three real GW events detected by LIGO during O3, showcasing its ability to generalize to GW signals without being explicitly trained on these waveforms.
This paper is structured as follows; in Section II, we discuss the state-of-the-art in glitch modelling and mitigation in LIGO and Virgo and state-of-the-art machine learning research for signal and glitch reconstruction. In Section III, we discuss our method, which includes an overview of our framework, our model architecture, the generation of training data, and the experiments implemented to show DeepExtractor’s efficacy. In Section IV, we present the results of the experiments and Section V discusses the conclusions of our research.
@@ II.1 GW glitch mitigation
A variety of methods have been developed to model and subtract glitches from the strain data that leverage machine learning [47] and parametric [48, 49] models. However, up to the end of O3, BayesWave [46, 50, 34, 51] and gwsubtract [52, 30, 53] have been the only algorithms used for modeling and subtracting glitches in LIGO-Virgo-KAGRA data analyses [53]. While gwsubtract relies on strongly correlated auxiliary witness sensors to linearly subtract glitches from the strain data, BayesWave models glitches using the detector strain only.
BayesWave models the strain $h(t)$ in each detector as a linear combination: $h(t)=n(t)+s(t)+g(t)$, where $n(t)$ represents Gaussian noise, $s(t)$ an astrophysical signal, and $g(t)$ a glitch. Non-Gaussian components, including signals and glitches, are represented as sums of sine-Gaussian (Morlet-Gabor) wavelets (see Eq. $4$ of [46]).
BayesWave performs Bayesian inference to estimate the wavelet parameters and the total number of wavelets of a glitch using a Reversible-Jump Markov Chain Monte Carlo (RJMCMC) sampler, which allows for the flexibility of models with variable dimensions.
@@
Figure 1: An overview of DeepExtractor’s reconstruction approach. Two seconds of detector strain $h(t)$ is processed into magnitude and phase spectrogramns using an STFT. This is fed through our network, which maps the instance to the magnitude and phase spectrograms of the underlying background noise. We then use an inverse STFT to yield the corresponding time series $n(t)$. An estimation of the underlying signal or glitch can be simply calculated as $h(t)-n(t)$.
II.2 Deep denoisers for GW data analysis
@@
U-Net models [45] have been applied to denoise BBH signals using time-frequency representations [60]. Adapted from seismic signal separation studies [61], this approach leverages the real and imaginary parts of spectrograms to predict separate masks for signal and noise. These masks, applied to input spectrograms, allow reconstruction of the time-domain signal via inverse STFT, yielding strong performance on real astrophysical events.
@@ III.1.1 Overview of framework
DeepExtractor is a deep learning framework designed to reconstruct signals or glitches in GW data with power exceeding the underlying detector noise. An overview of the framework is illustrated in FIG. 1. Built using PyTorch’s deep learning library [63], DeepExtractor’s codebase is publicly accessible on GitLab [^1] https://git.ligo.org/tom.dooney/deepextractor.git. DeepExtractor is agnostic to the origin and morphology of excess power in the detector, whether it is a signal or a glitch. However, in this paper, we primarily focus on the application of glitch reconstruction.
The framework operates under the first two models of BayesWave described in Section II.1, where the data contains either an astrophysical signal (signal-only mode) or a glitch (glitch-only mode). Currently, DeepExtractor does not handle the third model assumption, where a glitch overlaps with an astrophysical signal (signal+glitch mode). While this capability is critical for a fully-functional glitch mitigation algorithm, this study focuses on the simpler signal-only and glitch-only cases as a foundational step. Potential approaches for addressing the signal+glitch scenario are discussed in Section V.
We follow the assumption that a glitch $g(t)$ interacts additively with the strain data in a real detector setting. Our goal is to map the input mixture $h(t)$ to the noise component $n(t)$, where $h(t)$ is represented as:
@@ III.1.2 Training strategy
The DeepExtractor framework utilizes a neural network that directly outputs the background noise component from the data in the spectrogram domain. Our core idea is that the distribution of whitened background, which is approximately Gaussian stationary, is easier to model by our network compared to modelling the diverse classes of glitches.
@@
Neural networks can learn in either the time or time-frequency domain, each with trade-offs [64]. DeepExtractor is trained in the time-frequency domain using invertible STFTs, which provide rich signal representations that enhance learning [65, 66, 67, 68]. While many STFT-based models use only the magnitude spectrogram—assuming identical mixture (input) and source (target) phases [69]—recent work shows that incorporating phase via complex spectrograms improves performance [70, 71, 72]. Accordingly, we model both magnitude and phase components of the complex STFT spectrograms to enable accurate time-domain reconstruction of signals and glitches.
To provide the inputs and targets, we first transform the time-domain signals $h(t)$ and $n(t)$ into the time-frequency domain by computing their complex spectrograms, $h(t,f)$ and $n(t,f)$, using the STFT (see Section III.2.1). The spectrograms are then decomposed into magnitude and phase components, each with dimension $\mathbb{R}^{h\times w}$. These are then processed in two separate input and output channels of the network, using a data structure of dimension $\mathbb{R}^{2\times h\times w}$ (see FIG. 2). Thus, DeepExtractor takes the magnitude and phase components of $h(t,f)$ as input and outputs the corresponding components of the estimated noise $\hat{n}(t,f)$.
@@
All models in this study were trained on an NVIDIA A100 GPU (80GB memory). Our best-performing DeepExtractor configuration trained on the simulated dataset (see Section III.2.1) in approximately 8 hours, converging after 24 epochs. The subsequent transfer learning phase on real LIGO data (see Section III.2.2) required only 3 additional epochs to converge and completed in under 20 minutes.
@@ III.1.3 U-Net Architecture
Figure 2: The U-Net architecture featured in DeepExtractor applied to batches (bs) of STFT data ($\mathbb{R}^{2\times h\times w}$), illustrating its characteristic ‘U’-shaped structure. The network processes both the magnitude and phase components of the STFT simultaneously through two input and output channels.
DeepExtractor employs a U-Net architecture [45] for its ability to capture both high- and low-level features through skip connections. These connections concatenate encoder features with their corresponding decoder features (FIG. 2), preserving fine-grained details while supporting a large receptive field. This design has made U-Nets a state-of-the-art choice for signal denoising [74], acoustic source separation [75, 76], and image denoising [77]. U-Nets are particularly effective for spectrogram-based representations of signals and glitches spanning broad time-frequency ranges. Unlike deep autoencoders, which may lose information through aggressive compression, U-Nets preserve structural details throughout the network, enabling more accurate and faithful reconstructions [78].
DeepExtractor is a 2D model designed to process magnitude and phase spectrograms. It accepts two input channels and outputs two spectrograms of the same dimensionality, as illustrated in FIG. 2.
@@
Time-Domain Variant: We also experiment using a 1D time-domain counterpart of DeepExtractor’s U-Net architecture (UNET1D). This replaces all 2D operations with 1D equivalents and processes a single input/output channel. Specifically, the 1D model maps the raw, whitened time series $h(t)$ to the noise component $n(t)$, omitting the STFT and iSTFT transformations used in the 2D variant, as shown in FIG. 1.
@@ III.2.1 Simulated data
To properly validate our model, we first implement a fully simulated approach for both training and validation, simulating the background noise and glitches. We simulate white noise training targets $n(t)$ with a duration of 2 seconds at a sampling rate of $f_{s}=4,096\,$ Hz, yielding 8,092 data points. We generate 250,000 background samples for training and 25,000 for validation.
To create the corresponding input samples, we inject various analytical waveforms, represented by $g(t)$, into the background noise $n(t)$, as per Eq. 1. While BayesWave models glitches exclusively as linear combinations of sine-Gaussian wavelets, our approach broadens the space of basis waveforms to improve the network’s generalization capability. Specifically, we draw from five proxy glitch classes—chirp, sine, sine-Gaussian, Gaussian pulse, and ringdown—illustrated in FIG. 3.
The chirp class was selected for its resemblance to the frequency evolution characteristic of CBC signals. The sine class captures narrow-band features across varying frequencies. The sine-Gaussian class, inspired by BayesWave, along with the Gaussian pulse and ringdown classes, serves as an analytical proxy for gravitational-wave signals expected from a variety of burst sources [81]. Since these waveforms are commonly used in unmodelled burst searches, we investigate their suitability for unmodelled glitch reconstruction in our DeepExtractor framework.
Each waveform class has its own parameters that are randomly sampled during generation, such as frequency range (from 111\,1 Hz to the Nyquist frequency of $2,048\,$Hz), bandwidth, and duration (randomly chosen between 0.1250.1250.125\,0.125s and 222\,2s). The lower bound of 0.1250.1250.125\,0.125s was chosen since over 90% of high-confidence Gravity Spy glitches exceed this duration. Additionally, the analytical waveforms used for training often contain features—such as individual sine-Gaussian cycles—that evolve on timescales shorter than 0.1250.1250.125\,0.125s. While shorter durations could be considered, this threshold is expected to capture the majority of short-duration structures relevant for glitch reconstruction. For more information on these training classes, please see Appendix VII.1 and our code repository. The SNR is varied between 1 and 250, scaled according to
@@
where $g(f)$ and $S_{n}(f)$ represent the Fourier transform of the injected waveform and the detector noise PSD, respectively [82]. We set $S_{n}(f)$ to unity for convenience, since we are working with simulated whitened detector noise with a flat PSD.
To improve generalization and avoid overfitting, we inject linear combinations of glitches into each background sample, where:
@@
and $k\in[1,30]$. Futhermore, each $g_{i}(t)$, with $i\in[1,k]$, is randomly time-shifted within the 222\,2s duration. This procedure produces a diverse set of composite proxy glitches, which may or may not overlap in time and frequency. In the context of the training data, $g(t)$ in Eq. 1 can be replaced with $\mathbf{g}$ to reflect this linear combination. Inspired by BayesWave’s approach of modeling glitches as sums of sine-Gaussian wavelets, our setup injects between 1 and 30 glitches per sample, drawn from a broader five-class proxy glitch space.
Additionally, glitches can be scaled to low SNRs, so some training samples include injections with very little (or even negligible) contribution to the data. We hypothesize that this approach helps the model to learn features inherent to the background noise when glitches are not present in the data. We test this hypothesis on both simulated (Section III.3.1) and real detector noise (Section III.3.3).
To prepare the data for training DeepExtractor, we scale (standardize) the time-domain inputs and outputs using StandardScaler from sklearn [83], which transforms the data so that it has a mean of 0 and a standard deviation of 1. This standardization method assumes a Gaussian distribution of the data and is appropriate given the Gaussian-like characteristics of the background noise.
We fit the scaler exclusively to the model input (i.e., the glitch-injected samples) and use this input-fitted scaler to transform both the input and the background noise target. This ensures that both inputs and outputs occupy the same standardized space. While this data must be then transformed using the STFT to train DeepExtractor, it is this preprocessed data that is used to train the time-domain benchmarks (eg. UNET1D).
Short time Fourier transform (STFT): As mentioned in Section III.1.2, DeepExtractor and the 2D benchmarks are trained on spectrograms calculated from the above preprocessed data using the STFT. Since the time-domain data is already standardized, no additional scaling is performed on the spectrograms. FIG. 4 shows examples of the magnitude and phase components for both the input and target spectrograms.
We choose STFT parameters to ensure compliance with the Constant Overlap-Add (COLA) condition [84] (please see Appendix VII.2 for more information). A Hann window is employed to minimize spectral leakage, further improving the quality of the spectrogram representation [85]. These chosen settings yield complex spectrograms with dimensions of $257\times 257$, which we separate into magnitude and phase spectrograms. To investigate the effect of varying time and frequency resolutions on signal and glitch reconstruction, we also experiment with alternate parameter configurations that produce spectrograms of dimensions $129\times 129$ and $65\times 129$ while ensuring that all configurations satisfy the COLA condition.